1. Suppose the demand equation for computers by Teetan Ltd for the year 2017 is given by Qd= 1200-P and the supply equation is given by Qs= 120+3P. Find equilibrium price and analyse what would be the excess demand or supply if price changes to Rs 400 and Rs 120.

2.a. A business firms sells a good at the price of Rs 450.The firm has decided to reduce the price of good to Rs 350.Consequently, the quantity demanded for the good rose from 25,000 units to 35,000 units. Calculate the price elasticity of demand.

2.b. “There is a high cross elasticity of demand between new and old cars”. Discuss the statement by explaining the features of cross elasticity of demand. Also compare and contrast cross elasticity with other types of elasticities of demand.


1
Expert's answer
2020-11-30T16:38:04-0500

1) We are given:


Qd = 1,200 - P

Qs = 120 + 3P


At equilibrium, Qd = Qs

=> 1,200 - P = 120 + 3P

=> 1,200 - 120 = 3P + P

=> 1,080 = 4P

Dividing both sides by 4 gives,

P = $270


For equilibrium quantity, we substitute $270 in either Qd or Qs. Now, substituting $270 for P in Qs gives:

Qe = 120 + 3(270)

= 120 + 810

= 930 units.


Therefore, the equilibrium price is $270 and the equilibrium quantity is 930 units.



When price is $400:

Qd = 1,200 - 400

= 800 units


Qs = 120 + 3(400)

= 120 + 1200

= 1,320 units


Excess supply = Qs - Qd

= 1,320 - 800

= 520 units


Therefore, when price is $400, Qd exceeds Qs, generating an excess supply of 520 units


When price is $120:

Qd = 1200 - 120

1,080 units


Qs = 120 + 3(120)

= 480 units


Excess demand = Qd - Qs

= 1,080 - 480

= 600 units


Therefore, when price is $120, there is an excess supply of 600 units.



2 (a) Price elasticity of demand, PED:


"PED = \dfrac {\% \space ∆Quantity}{\% \space ∆Price}"


"\% \space ∆Q = \dfrac {35,000-25,000}{25,000}×100\%"


"= \dfrac {10,000}{25,000}×100\%"


"= 40\%"



"\% \space ∆P = \dfrac {350-450}{450} ×100\%"


"= -\dfrac {100}{450}×100\%"


"= -22.222222\%"


"\therefore \space PED = \dfrac {40\%}{-22.22222\%}"


"= -0.55555555"

"= \bold {-0.56}"



b) Cross elasticity of demand measures the responsiveness of quantity demanded of one good to changes in the price of another good. It is negative when goods are complements and positive when goods are substitutes. New cars are Substitutes of old cars. Therefore, the cross elasticity of demand between new and old cars is high and positive to imply that the two goods are strong substitutes.


On the contrary, price elasticity of demand measures the responsiveness of quantity demanded due to changes in the own price of the good. Income elasticity of demand measures the responsivess of quantity demanded to changes in disposable income. It is positive when goods are normal and negative when goods are inferior.